# Deep Equilibrium Convolutional Sparse Coding for Hyperspectral Image Denoising

### v1

@@ Abstract
Index Terms:
Hyperspectral image denoising, convolutional sparse coding, deep equilibrium model.
@@ I Introduction
Hyperspectral image (HSI), with its rich spectral information, is widely used in remote sensing [1], medical diagnostics [2], agriculture [3], and image recognition [4]. Increasing spectral resolution comes at the cost of reducing the number of photons received in each channel. Combined with atmospheric interference, this inevitably introduces noise during the sensing process, which can degrade subsequent applications. Therefore, HSI denoising is a crucial preprocessing step to ensure image quality and enable reliable downstream applications.
@@
The inherent spatial and spectral redundancy in HSIs allows clean HSIs to be represented as the combinations of few elements from a dictionary, which has driven the success of sparse coding in HSI denoising. Due to the high dimentionarity of reshaping the whole image into a vector, traditional sparse coding model typically requires partitioning the whole HSIs into multiple overlapping patches, thereby ignoring the shift-invariant properties of the data. As an alternative, convolutional sparse coding (CSC) employs a set of convolutional atoms to represent the image, naturally preserving the spatial relationships between pixels. Owing to this advantage, CSC has been widely applied to various inverse problems [5, 6, 7]. In the context of HSI denoising, Xiong et al. [8] and Yin et al. [9] extended 2D CSC models to 3D ones by employing 3D convolutions to jointly model local spatial-spectral correlations but fail to capture the global spectral correlations among bands. More recently, Tu et al. [10] enforced shared convolutional sparse coefficients across bands to capture inter-band spatial structural consistency. Nevertheless, this approach overlooks the local spatial-spectral correlations.
@@
Additionally, the CSC model is based on the hand-crafted sparsity prior and cannot enjoy the data-driven learning from data. To address this limitation, the deep unfolding technique have been used in [8, 10] to transform the iterative optimization process into a deep neural network with a fixed number of weight-tied layers, where each layer mimics one step of the original optimization. One major limitation of the deep unfolding models is the hardware memory constraints: during training, all intermediate activations must be stored for backpropagation, restricting the model depth and complicating the training. As a result, such unfolding networks are typically constrained to a small number of layers. However, the fixed number of iterations may not guarantee convergence, and running additional iterations at test time can lead to significant performance degradation [11, 12].
@@
In contrast, deep equilibrium models (DEQs) [13] offer a paradigm shift by treating the hidden layers of deep networks as the solution to a fixed-point problem. Instead of unfolding the optimization process into predefined depth, DEQs model infinite-depth networks and directly solve for the equilibrium (fixed) point during the forward pass. This fixed-point formulation makes DEQs particularly suitable for tasks involving iterative refinement, such as those frequently encountered in convex and non-convex optimization for model-driven methods. Additionally, DEQs leverage existing numerical solvers and implicit differentiation for forward evaluation and backward propagation, significantly reducing memory usage and improving training stability. In a nutshell, DEQ offers a manner to transform the traditional model-driven methods into a learnable architecture with convergence guarantees and preferable interpretability to bypass the limitation of hand-crafted priors. In the literature, Gilton et al. [11] demonstrated the improved effectiveness of DEQs in a variety of image reconstruction tasks over deep unfolding networks. Zhao et al.[14] extended the DEQ framework to recurrent neural networks (RNNs) and and Plug-and-Play (PnP) algorithms for snapshot compressive imaging. More recently, Geng et al.[15] applied DEQ to accelerate sampling in diffusion models.
@@ II-A HSI Denoising
In recent years, the paradigm of HSI denoising has shifted from model-driven to data-driven methods, with hybrid-driven approaches emerging as a promising direction that integrates the advantages of both. Defining an effective prior is crucial in traditional model-based methods. The strong spatial and spectral correlations in HSIs enable their representation in a low-dimensional subspace, underpinning the success of low-rankness and sparsity priors. Methods based on low-rank matrix or tensor decomposition [16, 17] and on rank minimization [18, 19, 20] denoise HSIs by extracting these low-rank structures. Meanwhile, the intrinsic redundancy of HSIs supports sparse representations over a few atoms from a dictionary. Peng et al. [21] reconstructed clean HSIs using either predefined or learned dictionaries, and Zhao et al. [22] employed sparse coding to capture both global spatial and local spectral redundancies. Furthermore, Zhuang et al. [23] preserved "rare pixels" containing critical information through collaborative sparsity. Total variation(TV) regularization–another form of transformed sparsity–has also proven effective to promote piecewise smoothness [24, 25]. Despite their interpretability and effectiveness, these model-driven approaches often demand extensive parameter tuning and entail high computational costs.
@@
Data-driven methods utilize various deep nerual network (DNN) architectures to model the intrinsic structure of HSI from data. Convolutional neural networks (CNNs) [26, 27, 28] can only model the local spatial-spectral correlation with limited receptive fields, restricting their ability to capture long-range dependencies that always important in image processing. Transformers treat non-overlapping image patches as sequences and leverage attention mechanisms to model long-range dependencies [29]. Compared to RGB images, HSIs possess significantly richer spectral resolution and exhibit unique spatial-spectral characteristics. Consequently, transformer-based methods for HSI denoising often emphasize the exploration of both spatial and spectral self-similarity. For instance, Peng et al. [21] employed a dual-branch network combining CNNs and transformers to separately capture spectral correlations and nonlocal spatial self-similarity. Li et al. [30] adopted non-local spatial self-attention and global spectral attention to fully exploit the inherent similarities along spatial and spectral dimensions. More recently, some approaches have combined self-attention with carefully designed low-rank modules to simultaneously capture spatial self-similarity and spectral low-rank structure [31, 32, 33]. However, the quadratic complexity of transformers in handling long sequences leads to substantial computational overhead, limiting their scalability and practicality in HSI applications. To address this, recent works have explored Mamba [34], a state space model architecture, for HSI denoising [35]. Data-driven nerual networks often require extensive empirical tuning and trial-and-error architecture design and can not share the well-established interpretability of model-based methods.
@@
Hybrid-driven methods have emerged as a promising research direction by combining the interpretability of model-driven approaches with the strong representation capabilities of data-driven techniques. To incorporate low-rank priors in HSIs, several studies [36, 37, 38, 39] have projected HSIs into low-dimensional spectral subspaces to facilitate efficient restoration. Similarly, sparsity priors have inspired the design of network architectures based on deep unfolding [8, 40, 41, 42, 43], where the network structure is derived by unfolding the optimization process of a model. For instance, Xiong et al. [44] converted the iterative soft-shrinkage algorithm of a multitask sparse representation model into an explainable network for enhanced HSI denoising. In [45], a total variation-based denoising model was unfolded into a learnable network. By further integrating a statistical feature injection module and a multiscale degradation guidance module, the method effectively recovers real structural details. Despite these advantages, deep unfolding networks often suffer from high memory consumption, instability and umerical issues arising in backpropagation, especially as the number of iterations increases, negatively impacting the reconstruction performance.
@@ II-B Deep Equilibrium Model
Traditional neural networks typically enhance their representational capacity by increasing the number of explicit layers. However, this strategy often leads to increased memory consumption and computational overhead. Recent research has shown that comparable or even superior performance can be achieved by sharing weights across all layers [46, 47]. Motivated by this insight, Bai et al. [13] introduced the DEQ, which abandons the conventional notion of a finite stack of layers. Instead, DEQ defines an implicit infinite-depth network by formulating a set of analytical conditions whose solution corresponds to the network’s output. Rather than unrolling layers, DEQ directly solves for the equilibrium point where the hidden representation becomes stable. This point can be found using efficient black-box root-finding methods such as Broyden’s method or Anderson acceleration. Crucially, DEQ leverages implicit differentiation to compute gradients during backpropagation without storing intermediate activations, enabling memory usage to remain constant at $\mathcal{O}(1)$. To enhance training stability, several follow-up studies have introduced techniques such as convergence-enforcing layers [48], Jacobian regularization [49], and phantom gradients [50]. Notably, the iterative optimization process in model-based methods naturally aligns with the equilibrium-seeking principle of DEQ, making it well-suited for tasks in imaging [11, 14] and denoising [51]. Beyond optimization, the multiscale fusion can also be framed within an equilibrium formulation. For instance, Bai et al. [52] proposed the multiscale deep equilibrium model (MDEQ), where inputs are injected at the highest resolution and propagated implicitly across scales to satisfy a joint equilibrium condition. In summary, the DEQ model provides an attractive manner to build the connection between the fixed-point optimization and learning-based techniques to enjoy both benefits. For this end, we adopt the DEQ model to transform the optimization of the CSC model for HSI denoising.
@@ III-B1 Weight-tied Convolutional Sparse Coding Layers
Eq. (13) is a weight-tied architecture with residual connections from the input to each layer until convergence. In other words, each layer incrementally refines the output by building upon the updates from the previous iteration. As the depth increases, the magnitude of these updates gradually diminishes, leading to a regime of diminishing returns, where additional layers contribute progressively less until the network reaches a stable equilibrium. As pointed in [13], such a design offers several advantages. First, weight sharing serves as a form of implicit regularization that stabilizes training and promotes generalization. Second, it significantly reduces the number of trainable parameters, resulting in more compact models. Third, unlike deep unfolding models that require a pre-defined, fixed number of layers, this iterative structure can theoretically be unrolled to arbitrary depth, aligning naturally with the principles of fixed-point optimization. This architecture seamlessly integrates with numerical solvers to reach a stable equilibrium and therefore has convergence guarantee.
@@ III-B2 Forward and Backward Pass
Since the forward pass of DEQ does not rely on explicit iterations, gradient computation cannot be automatically performed with Pytorch or Tensorflow, during the backward propagation process. According to the derivation by Bai et al. [13], the gradient with respect to the network parameters be learned with constant memory using implicit differentiation and is calculated as
@@
Since the exact computation of the Jacobian-inverse $\left(\mathbf{I}-\frac{\partial f_{\Theta}(\alpha^{*},\mathbf{z})}{\partial\alpha^{*}}\right)^{-1}$ requires high computational costs, we approximates it using phantom gradient [50]. Specifically, consider the Neumann series expansion of the Jacobian-inverse:
@@ III-C3 The Detail Enhancement Module
where $\mathrm{Conv}_{1}$ is the $3\times 3\times 3$ 3D convolution and DConv denotes the difference convolution operator [53]. Due to the fact that conventional CNNs are optimized from random initialization, they often struggle to effectively focus on extracting image edges and fine details during training. In contrast, traditional edge detection operators characterize abrupt intensity changes and fine structural features by leveraging differential information from the image. Therefore, DConv[53] combines conventional edge detection operators with CNNs, enabling more precise capture of image gradient information while fully exploiting the powerful representational capability of deep learning. As illustrated in Fig.3, we extend conventional edge detection operators to design a DConv module tailored for HSI. It consists of a standard 3D convolution for extracting intensity information, along with four edge extraction operators specifically designed for HSI to capture gradient information. In particular, the central difference operator enhances image sharpness, the inter-band difference operator captures variations along the spectral dimension, and the horizontal difference and vertical difference operators in the spatial domain extract horizontal and vertical edge information, respectively. This intermediate output is then passed through two cascaded 3D convolutional layers, which serve as an attention that adaptively emphasizes regions rich in image detail information.
@@ IV-A1 Benchmarked Models
The comparison involves 12 methods, comprising 7 model-driven methods, 3 data-driven methods, and 2 hybrid-driven methods. The model-driven methods include BM4D [54], MTSNMF [55], LLRT [56], NGMeet [17], LRMR [57], E-3DTV [25], and 3DlogTNN [18]. The data-driven methods include SST [30], TRQ3D [58] and SERT [31], while the hybrid-driven methods based on sparse priors include T3SC [41] and MTSNN++ [59].
@@ IV-A3 Noise Patterns
Non-independent and identically distributed (non-i.i.d.) Gaussian noise and mixture noise were considered in the synthetic experiment. In addition, we adopted the "noise with spectrally correlated variance" model proposed by Bodrito et al. [41]. The detailed noise settings are described as follows:
@@
[54] [55] [56] [17] [57] [25] [18] [30] [58] [31] [41] [59] (Ours)
@@
Figure 4: Denoising results on the Nachal_08231214Nachal\_0823-1214 HSI from the ICVL dataset under the non-i.i.d. Gaussian noise with $\sigma\in[0,95]$. The false-color images are generated by combining bands 31, 17, and 2. (a) Clean. (b) Noisy. (c) BM4D [54]. (d) MTSNMF [55]. (e) LLRT [56]. (f) NGMeet [17]. (g) LRMR [57]. (h) E-3DTV [25]. (i) 3DlogTNN [18]. (j) SST [30]. (k) TRQ3D [58]. (l) SERT [31] (m) T3SC [41]. (n) MTSNN++ [59]. (o) DECSC.

### v2

@@ I Introduction
Hyperspectral image (HSI), with its rich spectral information, is widely used in remote sensing [cloutis1996review], medical diagnostics [fei2019hyperspectral], agriculture [kersting2012pre], and image recognition [fauvel2012advances]. Increasing spectral resolution comes at the cost of reducing the number of photons received in each channel. Combined with atmospheric interference, this inevitably introduces noise during the sensing process, which can degrade subsequent applications. Therefore, HSI denoising is a crucial preprocessing step to ensure image quality and enable reliable downstream applications.
@@
The inherent spatial and spectral redundancy in HSIs allows clean HSIs to be represented as the combinations of few elements from a dictionary, which has driven the success of sparse coding in HSI denoising. Due to the high dimentionarity of reshaping the whole image into a vector, traditional sparse coding model typically requires partitioning the whole HSIs into multiple overlapping patches, thereby ignoring the shift-invariant properties of the data. As an alternative, convolutional sparse coding (CSC) employs a set of convolutional atoms to represent the image, naturally preserving the spatial relationships between pixels. Owing to this advantage, CSC has been widely applied to various inverse problems [gu2015convolutional, bao2019convolutional, hu2017convolutional]. In the context of HSI denoising, Xiong et al. [xiong2022spatial] and Yin et al. [yin2024deep] extended 2D CSC models to 3D ones by employing 3D convolutions to jointly model local spatial-spectral correlations but fail to capture the global spectral correlations among bands. More recently, Tu et al. [Tu2024] enforced shared convolutional sparse coefficients across bands to capture inter-band spatial structural consistency. Nevertheless, this approach overlooks the local spatial-spectral correlations.
@@
Additionally, the CSC model is based on the hand-crafted sparsity prior and cannot enjoy the data-driven learning from data. To address this limitation, the deep unfolding technique have been used in [xiong2022spatial, Tu2024] to transform the iterative optimization process into a deep neural network with a fixed number of weight-tied layers, where each layer mimics one step of the original optimization. One major limitation of the deep unfolding models is the hardware memory constraints: during training, all intermediate activations must be stored for backpropagation, restricting the model depth and complicating the training. As a result, such unfolding networks are typically constrained to a small number of layers. However, the fixed number of iterations may not guarantee convergence, and running additional iterations at test time can lead to significant performance degradation [gilton2021deep, Monga2021].
@@
In contrast, deep equilibrium models (DEQs) [bai2019deep] offer a paradigm shift by treating the hidden layers of deep networks as the solution to a fixed-point problem. Instead of unfolding the optimization process into predefined depth, DEQs model infinite-depth networks and directly solve for the equilibrium (fixed) point during the forward pass. This fixed-point formulation makes DEQs particularly suitable for tasks involving iterative refinement, such as those frequently encountered in convex and non-convex optimization for model-driven methods. Additionally, DEQs leverage existing numerical solvers and implicit differentiation for forward evaluation and backward propagation, significantly reducing memory usage and improving training stability. In a nutshell, DEQ offers a manner to transform the traditional model-driven methods into a learnable architecture with convergence guarantees and preferable interpretability to bypass the limitation of hand-crafted priors. In the literature, Gilton et al. [gilton2021deep] demonstrated the improved effectiveness of DEQs in a variety of image reconstruction tasks over deep unfolding networks. Zhao et al.[zhao2023deep] extended the DEQ framework to recurrent neural networks (RNNs) and and Plug-and-Play (PnP) algorithms for snapshot compressive imaging. More recently, Geng et al.[geng2023one] applied DEQ to accelerate sampling in diffusion models.
@@ II-A HSI Denoising
In recent years, the paradigm of HSI denoising has shifted from model-driven to data-driven methods, with hybrid-driven approaches emerging as a promising direction that integrates the advantages of both. Defining an effective prior is crucial in traditional model-based methods. The strong spatial and spectral correlations in HSIs enable their representation in a low-dimensional subspace, underpinning the success of low-rankness and sparsity priors. Methods based on low-rank matrix or tensor decomposition [zheng2020double, He2022NGMeet] and on rank minimization [Zheng20203DlogTNN, xie2021adaptive, chen2022hyperspectral] denoise HSIs by extracting these low-rank structures. Meanwhile, the intrinsic redundancy of HSIs supports sparse representations over a few atoms from a dictionary. Peng et al. [peng2014decomposable] reconstructed clean HSIs using either predefined or learned dictionaries, and Zhao et al. [Zhao2015Hyperspectral] employed sparse coding to capture both global spatial and local spectral redundancies. Furthermore, Zhuang et al. [zhuang2020hyperspectral] preserved "rare pixels" containing critical information through collaborative sparsity. Total variation(TV) regularization–another form of transformed sparsity–has also proven effective to promote piecewise smoothness [wang2017hyperspectral, Peng2020E-3DTV]. Despite their interpretability and effectiveness, these model-driven approaches often demand extensive parameter tuning and entail high computational costs.
@@
Data-driven methods utilize various deep nerual network (DNN) architectures to model the intrinsic structure of HSI from data. Convolutional neural networks (CNNs) [yuan2018hyperspectral, chang2018hsi, zhang2019hybrid] can only model the local spatial-spectral correlation with limited receptive fields, restricting their ability to capture long-range dependencies that always important in image processing. Transformers treat non-overlapping image patches as sequences and leverage attention mechanisms to model long-range dependencies [dosovitskiy2020image]. Compared to RGB images, HSIs possess significantly richer spectral resolution and exhibit unique spatial-spectral characteristics. Consequently, transformer-based methods for HSI denoising often emphasize the exploration of both spatial and spectral self-similarity. For instance, Peng et al. [peng2014decomposable] employed a dual-branch network combining CNNs and transformers to separately capture spectral correlations and nonlocal spatial self-similarity. Li et al. [li2022spatialspectral] adopted non-local spatial self-attention and global spectral attention to fully exploit the inherent similarities along spatial and spectral dimensions. More recently, some approaches have combined self-attention with carefully designed low-rank modules to simultaneously capture spatial self-similarity and spectral low-rank structure [li2023spectral, li2024latent, tan2024low]. However, the quadratic complexity of transformers in handling long sequences leads to substantial computational overhead, limiting their scalability and practicality in HSI applications. To address this, recent works have explored Mamba [Gu2023], a state space model architecture, for HSI denoising [fu2024ssumamba]. Data-driven nerual networks often require extensive empirical tuning and trial-and-error architecture design and can not share the well-established interpretability of model-based methods.
@@
Hybrid-driven methods have emerged as a promising research direction by combining the interpretability of model-driven approaches with the strong representation capabilities of data-driven techniques. To incorporate low-rank priors in HSIs, several studies [Xiong2022MAC-Net, zhang2023combined, Chen2024, zhuang2023eigenimage2eigenimage] have projected HSIs into low-dimensional spectral subspaces to facilitate efficient restoration. Similarly, sparsity priors have inspired the design of network architectures based on deep unfolding [xiong2022spatial, Zeng2025, bodrito2021T3SC, xiong2022smds, li2023pixel], where the network structure is derived by unfolding the optimization process of a model. For instance, Xiong et al. [Xiong2023] converted the iterative soft-shrinkage algorithm of a multitask sparse representation model into an explainable network for enhanced HSI denoising. In [Li2023], a total variation-based denoising model was unfolded into a learnable network. By further integrating a statistical feature injection module and a multiscale degradation guidance module, the method effectively recovers real structural details. Despite these advantages, deep unfolding networks often suffer from high memory consumption, instability and umerical issues arising in backpropagation, especially as the number of iterations increases, negatively impacting the reconstruction performance.
@@ II-B Deep Equilibrium Model
Traditional neural networks typically enhance their representational capacity by increasing the number of explicit layers. However, this strategy often leads to increased memory consumption and computational overhead. Recent research has shown that comparable or even superior performance can be achieved by sharing weights across all layers [bai2018trellis, dehghani2018universal]. Motivated by this insight, Bai et al. [bai2019deep] introduced the DEQ, which abandons the conventional notion of a finite stack of layers. Instead, DEQ defines an implicit infinite-depth network by formulating a set of analytical conditions whose solution corresponds to the network’s output. Rather than unrolling layers, DEQ directly solves for the equilibrium point where the hidden representation becomes stable. This point can be found using efficient black-box root-finding methods such as Broyden’s method or Anderson acceleration. Crucially, DEQ leverages implicit differentiation to compute gradients during backpropagation without storing intermediate activations, enabling memory usage to remain constant at $\mathcal{O}(1)$. To enhance training stability, several follow-up studies have introduced techniques such as convergence-enforcing layers [winston2020monotone], Jacobian regularization [bai2021stabilizing], and phantom gradients [geng2021training]. Notably, the iterative optimization process in model-based methods naturally aligns with the equilibrium-seeking principle of DEQ, making it well-suited for tasks in imaging [gilton2021deep, zhao2023deep] and denoising [gkillas2023connections]. Beyond optimization, the multiscale fusion can also be framed within an equilibrium formulation. For instance, Bai et al. [bai2020multiscale] proposed the multiscale deep equilibrium model (MDEQ), where inputs are injected at the highest resolution and propagated implicitly across scales to satisfy a joint equilibrium condition. In summary, the DEQ model provides an attractive manner to build the connection between the fixed-point optimization and learning-based techniques to enjoy both benefits. For this end, we adopt the DEQ model to transform the optimization of the CSC model for HSI denoising.
@@ III-B1 Weight-tied Convolutional Sparse Coding Layers
Eq. (13) is a weight-tied architecture with residual connections from the input to each layer until convergence. In other words, each layer incrementally refines the output by building upon the updates from the previous iteration. As the depth increases, the magnitude of these updates gradually diminishes, leading to a regime of diminishing returns, where additional layers contribute progressively less until the network reaches a stable equilibrium. As pointed in [bai2019deep], such a design offers several advantages. First, weight sharing serves as a form of implicit regularization that stabilizes training and promotes generalization. Second, it significantly reduces the number of trainable parameters, resulting in more compact models. Third, unlike deep unfolding models that require a pre-defined, fixed number of layers, this iterative structure can theoretically be unrolled to arbitrary depth, aligning naturally with the principles of fixed-point optimization. This architecture seamlessly integrates with numerical solvers to reach a stable equilibrium and therefore has convergence guarantee.
@@ III-B2 Forward and Backward Pass
Since the forward pass of DEQ does not rely on explicit iterations, gradient computation cannot be automatically performed with Pytorch or Tensorflow, during the backward propagation process. According to the derivation by Bai et al. [bai2019deep], the gradient with respect to the network parameters be learned with constant memory using implicit differentiation and is calculated as
@@
Since the exact computation of the Jacobian-inverse $\left(\mathbf{I}-\frac{\partial f_{\Theta}(\alpha^{*},\mathbf{z})}{\partial\alpha^{*}}\right)^{-1}$ requires high computational costs, we approximates it using phantom gradient [geng2021training]. Specifically, consider the Neumann series expansion of the Jacobian-inverse:
@@ III-C3 The Detail Enhancement Module
where $\mathrm{Conv}_{1}$ is the $3\times 3\times 3$ 3D convolution and DConv denotes the difference convolution operator [yu2020searching]. Due to the fact that conventional CNNs are optimized from random initialization, they often struggle to effectively focus on extracting image edges and fine details during training. In contrast, traditional edge detection operators characterize abrupt intensity changes and fine structural features by leveraging differential information from the image. Therefore, DConv[yu2020searching] combines conventional edge detection operators with CNNs, enabling more precise capture of image gradient information while fully exploiting the powerful representational capability of deep learning. As illustrated in Fig.3, we extend conventional edge detection operators to design a DConv module tailored for HSI. It consists of a standard 3D convolution for extracting intensity information, along with four edge extraction operators specifically designed for HSI to capture gradient information. In particular, the central difference operator enhances image sharpness, the inter-band difference operator captures variations along the spectral dimension, and the horizontal difference and vertical difference operators in the spatial domain extract horizontal and vertical edge information, respectively. This intermediate output is then passed through two cascaded 3D convolutional layers, which serve as an attention that adaptively emphasizes regions rich in image detail information.
@@ IV-A1 Benchmarked Models
The comparison involves 12 methods, comprising 7 model-driven methods, 3 data-driven methods, and 2 hybrid-driven methods. The model-driven methods include BM4D [Maggioni2013BM4D], MTSNMF [Ye2015MTSNMF], LLRT [Chang2017LLRT], NGMeet [He2022NGMeet], LRMR [Zhang2014LRMR], E-3DTV [Peng2020E-3DTV], and 3DlogTNN [Zheng20203DlogTNN]. The data-driven methods include SST [li2022spatialspectral], TRQ3D [Pang2022TRQ3DNet] and SERT [li2023spectral], while the hybrid-driven methods based on sparse priors include T3SC [bodrito2021T3SC] and MTSNN++ [xiong2023multitask].
@@ IV-A3 Noise Patterns
Non-independent and identically distributed (non-i.i.d.) Gaussian noise and mixture noise were considered in the synthetic experiment. In addition, we adopted the "noise with spectrally correlated variance" model proposed by Bodrito et al. [bodrito2021T3SC]. The detailed noise settings are described as follows:
@@
[Maggioni2013BM4D] [Ye2015MTSNMF] [Chang2017LLRT] [He2022NGMeet] [Zhang2014LRMR] [Peng2020E-3DTV] [Zheng20203DlogTNN] [li2022spatialspectral] [Pang2022TRQ3DNet] [li2023spectral] [bodrito2021T3SC] [xiong2023multitask] (Ours)
@@
Figure 4: Denoising results on the Nachal_08231214Nachal\_0823-1214 HSI from the ICVL dataset under the non-i.i.d. Gaussian noise with $\sigma\in[0,95]$. The false-color images are generated by combining bands 31, 17, and 2. (a) Clean. (b) Noisy. (c) BM4D [Maggioni2013BM4D]. (d) MTSNMF [Ye2015MTSNMF]. (e) LLRT [Chang2017LLRT]. (f) NGMeet [He2022NGMeet]. (g) LRMR [Zhang2014LRMR]. (h) E-3DTV [Peng2020E-3DTV]. (i) 3DlogTNN [Zheng20203DlogTNN]. (j) SST [li2022spatialspectral]. (k) TRQ3D [Pang2022TRQ3DNet]. (l) SERT [li2023spectral] (m) T3SC [bodrito2021T3SC]. (n) MTSNN++ [xiong2023multitask]. (o) DECSC.